Dice Game Scorecard

UPPER SEC- TION

How to Score Game #1 Game #2 Game #3 Game #4 Game #5 Game #6

Ones  Sum of
Twos  Sum of
Threes  Sum of
Fours  Sum of
Fives  Sum of
Sixes  Sum of
TOTAL SCORE  ⟶

BONUS (if 63 or higher)

SCORE 35

TOTAL of Up- per Section

⟶

LOWER SEC- TION

{3 of a \atop Kind} Add Total of All Dice
{4 of a \atop Kind} Add Total of All Dice
Full House  25 Points

Small Straight

35 Points

Big Straight  40 Points
{5 of a \atop Kind}  50 Points
Chance  Total of all 5 Dice

BONUS!

TOTAL of Up- per Section

⟶

TOTAL of Lower Section

⟶
GRAND TO- TAL  ⟶

1